# Teaching Large Language Models to Regress Accurate Image Quality Scores Using Score Distribution

### v2

@@ 4.2 Score Estimation during Inference
Here we follow Q-Align [70] to apply a close-set softmax on the five levels to obtain the probabilities, $p_{i}^{pred}$, to prevent the influence of other textual tokens. Our statistics in Tab. A2 reveal that the trained MLLM will always predict the five levels, thus applying softmax on either the five levels or all textual tokens yields nearly identical results. As given in Tab. 5, our predicted score distribution, $\mathcal{N}(\mu^{pred},(\sigma^{pred})^{2})$, aligns well with human annotations.
@@ 4.3 Joint Training on Multiple IQA Datasets
As demonstrated in [87, 98], different IQA datasets exhibit distinct distributions. Due to this dataset variation, images from different datasets with nearly identical scores (i.e., linearly re-scaled scores) can have drastically different perceptual quality (see examples in Fig. A1). The reason is that different datasets have different reflections between quality scores and visual quality. This disparity makes it infeasible to merely regress absolute quality scores for multi-dataset co-training. To address this challenge, UNIQUE [87] applies fidelity loss [63] based on Thurstone’s model [60]. The fidelity loss is based on the principal that IQA emphasizes the ranking relationships within the same dataset, rather than the absolute values of quality scores. Specifically, two images are sampled from the same dataset, whose quality scores can be comparable. These two images are then compared using human-annotated quality scores. The fidelity loss encourages the model to learn such comparison results, which facilitates the model to focus more on the ranking relationships in each dataset, instead of the absolute values of quality scores. Thus, the fidelity loss helps multi-dataset co-training. Our soft label is naturally compatible with the fidelity loss, since our method can predict the score distribution, which is what the fidelity loss needs.
@@
Training with fidelity loss. Following [87], we choose fidelity loss as the similarity measure to train the model as:
@@
Handcrafted NIQE [41] 0.533 / 0.530 0.679 / 0.664 0.468 / 0.405 0.195 / 0.161 0.493 / 0.449 0.560 / 0.533 0.718 / 0.628 0.147 / 0.100
BRISQUE [40] 0.225 / 0.226 0.490 / 0.406 0.429 / 0.356 0.267 / 0.232 0.361 / 0.313 0.541 / 0.497 0.740 / 0.556 0.108 / 0.054
Non-MLLM Deep-learning NIMA [57] 0.896 / 0.859 0.838 / 0.856 0.532 / 0.535 0.390 / 0.399 0.814 / 0.771 0.715 / 0.654 0.695 / 0.649 0.561 / 0.467
HyperIQA [55] 0.917 / 0.906 0.791 / 0.788 0.506 / 0.468 0.410 / 0.403 0.772 / 0.749 0.702 / 0.640 0.752 / 0.717 0.485 / 0.383
DBCNN [86] 0.884 / 0.875 0.812 / 0.806 0.497 / 0.484 0.384 / 0.381 0.773 / 0.755 0.730 / 0.641 0.586 / 0.572 0.485 / 0.385
MUSIQ [26] 0.924 / 0.929 0.868 / 0.863 0.575 / 0.556 0.431 / 0.431 0.789 / 0.830 0.722 / 0.630 0.771 / 0.710 0.565 / 0.467
CLIP-IQA+ [64] 0.909 / 0.895 0.866 / 0.864 0.653 / 0.654 0.427 / 0.419 0.832 / 0.805 0.736 / 0.685 0.772 / 0.719 0.427 / 0.316
ManIQA [73] 0.849 / 0.834 0.768 / 0.758 0.499 / 0.465 0.457 / 0.452 0.849 / 0.832 0.723 / 0.636 0.623 / 0.627 0.512 / 0.401
MLLM-based Compare2Score [98] 0.923 / 0.910 0.867 / 0.860 0.500 / 0.453 0.354 / 0.342 0.786 / 0.772 0.777 / 0.671 0.735 / 0.705 0.474 / 0.413
Q-Align [70] 0.941 / 0.940 0.886 / 0.887 0.674 / 0.684 0.403 / 0.419 0.853 / 0.860 0.772 / 0.735 0.785 / 0.737 0.554 / 0.483
@@
0 KonIQ, SPAQ Q-Align [70] 0.943 / 0.940 0.933 / 0.931 0.692 / 0.708 0.401 / 0.411 0.883 / 0.879 0.795 / 0.727 0.671 / 0.568 0.795 / 0.767
@@
1 KonIQ, SPAQ, KADID Q-Align [70] 0.945 / 0.938 0.933 / 0.931 0.935 / 0.934 0.409 / 0.420 0.887 / 0.883 0.788 / 0.733 0.829 / 0.808 0.876 / 0.845
@@
2 KonIQ, SPAQ, KADID, PIPAL Q-Align [70] 0.926 / 0.932 0.917 / 0.920 0.950 / 0.954 0.702 / 0.671 0.853 / 0.845 0.765 / 0.722 0.811 / 0.795 0.838 / 0.789
@@ 5.1 Datasets, Baselines, Metrics, and Details
Datasets. For a comprehensive evaluation, we include a wide range of IQA datasets: (a) in-the-wild images: KonIQ [21], SPAQ [17], LIVE-Wild [18], and FLIVE [78]; (b) synthetic distortions: KADID [31], TID2013 [47], and CSIQ [29]; (c) model-processed distortions: PIPAL [24]; (d) AI-generated images: AGIQA-3K [30]. Among these datasets, KonIQ, SPAQ, KADID, and PIPAL are split into training and test sets. Other datasets are used to assess the out-of-distribution generalization ability. The Mean Opinion Scores (MOS) of these datasets are normalized to the range of [1,5], with the variances normalized accordingly.
Baseline methods. To demonstrate the effectiveness of our distribution-based soft label, we primarily compare with Q-Align [70], which utilizes the one-hot label. The base models for DeQA-Score and Q-Align are kept identical for fair comparison. We also include previous IQA methods as baselines, including handcrafted methods: NIQE [41] and BRISQUE [40], and deep-learning-based methods: NIMA [57], HyperIQA [55], DBCNN [86], MUSIQ [26], CLIP-IQA+ [64], ManIQA [73], and Compare2Score [98].
Metrics. Following [26, 70, 73], we use the Pearson Linear Correlation Coefficient (PLCC) and Spearman Rank-order Correlation Coefficient (SRCC) as metrics to evaluate the score regression performance. Additionally, we employ KL divergence, JS divergence, and Wasserstein distance to assess the results of score distribution prediction.
Implementation details. We use mPLUG-Owl2 [75] as our base model, with a CLIP-pretrained ViT-L [49] as the image encoder, a six-layer Q-Former as the visual abstractor, and LLaMA-2-7B [62] as the LLM. The number of visual tokens output by the vision abstractor is set to 64, and the hidden dimension of LLM is 4096. The pretrained weights of mPLUG-Owl2 are used for model initialization. The weighting term $\gamma$ is empirically set to 0.05, adjusting the two loss terms roughly to the same scale. We adopt AdamW [23] as the optimizer. The initial learning rate is set to 2e-5 and decays gradually using the cosine decay strategy. Our model is trained with a batch size of 64 for 3 epochs. Using 8 NVIDIA RTX A6000 GPUs, the training process on KonIQ dataset is completed in around 1.5 hours.
@@ 5.2 Results of Score Regression
Score regression results of single-dataset training. We first train our model on a single dataset, i.e., KonIQ, and then evaluate the model on both in-distribution and out-of-distribution datasets, following [28, 70]. The score regression results are shown in Tab. 3. On the in-distribution KonIQ test set, our model achieves superior score regression results compared to all baseline methods, e.g., a 1.3% improvement in PLCC over Q-Align. Furthermore, on out-of-distribution datasets, our DeQA-Score maintains its advantage with the best overall results across nearly all benchmarks, e.g., a 4.6% improvement in PLCC on LIVE-Wild. These findings confirm the effectiveness of our method.
@@ 5.3 Results of Score Distribution Prediction
Datasets Q-Align [70] DeQA-Score (Ours)
@@ 5.4 Ablation Studies
Effectiveness of soft label and fidelity loss is studied and shown in Tab. 6. In #1, “Uncertainty” is utilized to verify the effects of adding annotation variances without considering the score distortion. We directly append uncertainty within the responses, i.e., “The quality of this image is <<<level>>> with <<<uncertainty_level>>> uncertainty”. The <<<uncertainty_level>>> has five possible values: {“minimal”, “low”, “moderate”, “high”, “severe”}, derived by the one-hot discretization. In #2, model trained on one-hot label can not predict a reliable variance, thus we employ binary fidelity loss [89], which does not require predicted variance.
@@ B.1 Details of Dataset Construction
According to the $3\sigma$ rule, nearly all samples fall within the range $[\mu-3\sigma,\mu+3\sigma]$. Thus, if $3\sigma$ is less than half of the level interval, only the two nearest levels have non-zero probabilities. This criterion provides a way to define the threshold for small variance as $23\sigma\leq d/$ 2, which simplifies to $\sigma\leq 0.17$. To allow for a slightly relaxed threshold, the threshold for small (normalized) variance is set to $(0.2)^{2}$.
@@ B.2 Details of Methodology
Model architecture. DeQA-Score adopts the architecture of mPLUG-Owl2 [75], structured as follows. Specifically, the input images and the question texts are first tokenized, then fused, finally processed by the Large Language Model (LLM) for response generation. (a) Tokenizing input images and texts. We use a pre-trained CLIP pre-trained ViT-L/14 [49] as the image encoder to convert the input images into visual tokens, with each token having a channel of 1024. The texts are tokenized into textual tokens using the SentencePiece tokenizer [27], with each token having a channel of 4096. To bridge the different embedding spaces of visual and textual tokens, a trainable image abstractor, which is a six-layer transformer network, is implemented to map the vision tokens to the hidden dimension of the LLM, which is 4096. The abstractor can also significantly reduce the number of vision tokens to 64, relieving the computing pressure. (b) Token fusion. We integrate the visual tokens into pre-defined positions within the textual tokens as token fusion. (c) Response generation using LLM. The fused tokens are fed into an LLM, which is LLaMA-2-7B [62], to generate the final response. The LLM can be either LoRA-tuned [22] or fully tuned, and their results are similar when the vision components are trainable, as shown in Tab. A7.
@@ B.3 Details of Training and Inference
Training Setup. The pre-trained weights of mPLUG-Owl2 are used for model initialization. The loss weighting term is set to 0.05, bringing the two loss terms to roughly the same scale. We adopt AdamW [23] as the optimizer, with an initial learning rate of 2e-5 that gradually decays using a cosine decay strategy. A warmup strategy is applied with a warmup ratio of 0.03. Our model is trained with a batch size of 64 for 3 epochs. Both the vision encoder and abstractor are trainable, and the LLM is fully tuned unless specified. As shown in Tab. A7, LoRA-tuned LLM achieves comparable performance to fully-tuned LLM, providing an alternative for scenarios with limited computation resources. Using 8 RTX A6000 GPUs, training on the KonIQ dataset is completed in about 1.5 hours, while training on the KonIQ, SPAQ, and KADID datasets takes around 4 hours.
Explanation of closed-set softmax during inference. In the main paper, we follow Q-Align [70] by applying a closed-set softmax on the five levels to compute the probabilities $p_{i}^{\text{pred}}$, thereby avoiding the influence of other textual tokens. As shown in our statistics in Tab. A2, the probability sum of the five levels after training is very close to 1, indicating that the trained MLLM consistently predicts one of the five levels. Thus, applying softmax over the five levels or across all textual tokens yields nearly identical results.
@@ Appendix C More Results
The results are given in Tab. A3, where the KL divergence of Q-Align is quite large. We explain this as follows. As in Figs. 6, A4, A3 and A5, Q-Align tends to predict a single level token, bringing a quite small predicted variance. That means, the first variance, $\sigma_{1}^{2}$ in Eq. A3, is quite small ($\to 0$), thus the KL divergence goes extremely large ($\to$+$\infty$ + ∞).
@@
Results on more AI-generated images. In the main paper, we have included an IQA dataset with AI-generated images, AGIQA-3K [30], for evaluation. Here we further provide evaluation results on three additional AIGC datasets including AIGCIQA2023 [65], AGIN [11], and AGIQA-1K [90] in Tab. A4. Both Q-Align and our DeQA-Score are trained on the KonIQ, SPAQ, and KADID datasets, and then directly evaluated on these unseen AIGC datasets. Our DeQA-Score consistently outperforms the baseline method.
@@
LIQE [89] 0.907 / 0.922 0.916 / 0.921 0.928 / 0.929 0.503 / 0.493 0.855 / 0.822 0.689 / 0.650 0.881 / 0.854 0.792 / 0.794
Compare2Score [98] 0.941 / 0.929 0.929 / 0.927 0.952 / 0.949 0.446 / 0.440 0.868 / 0.856 0.787 / 0.733 0.836 / 0.809 0.879 / 0.830
Q-Align [70] 0.945 / 0.938 0.933 / 0.931 0.935 / 0.934 0.409 / 0.420 0.887 / 0.883 0.788 / 0.733 0.829 / 0.808 0.876 / 0.845
@@
Table A7: Ablation studies on training / fixing various model components with PLCC / SRCC metrics. “Enc.” means vision encoder, and “Abs.” represents vision abstractor. “LLM LoRA” or “LLM Full” represents the LLM is LoRA-tuned [22] or fully tuned. “✔” means this component is trained. The models are trained on the KonIQ, SPAQ, and KADID datasets. The results show that training vision encoder and abstractor significantly improves the performance. When vision encoder and abstractor are trained, fully-tuned LLM only shows a slight advantage over LoRA-tuned LLM.
@@
Table A8: Low-level perception results on Q-Bench [68] of co-training with or pre-training on score regression tasks. The results of #0, #1, and #2 are borrowed from [69]. Co-training with or pre-training on score regression tasks stably improves the performance.
@@
0 From the scratch [69] 0.7218 0.5796 0.5619 0.5668 0.6921 0.5329 0.7265 0.6161
1 High-level co-training [69] 0.7564 0.6704 0.5903 0.7101 0.6528 0.6316 0.6980 0.6756
2 High-level pre-training [69] 0.7600 0.6504 0.6166 0.6595 0.6875 0.6546 0.7388 0.6796
@@
Table A9: Co-training with Q-Instruct [69] leads to an obvious reduction in score regression performance. Exploring better strategies to combine language-based quality description tasks with score regression tasks is left as our future work.
@@
More baselines on multi-dataset training. In the main paper, we primarily compare with Q-Align. Here we add the comparison results with more baselines including LIQE [89] and Compare2Score [98] in Tab. A5. Our DeQA-Score consistently outperforms these baselines.
Ablation studies on level number are conducted in Tab. A7. The performance improves with moderately larger level numbers (i.e., 6$\sim$8), as discretization becomes more accurate. However, the performance declines with much larger level numbers (i.e., 10$\sim$12), due to the increased difficulty in level prediction (more classification categories).
Ablation studies on training / fixing components are summarized in Tab. A7. The vision encoder and abstractor can be either fixed or trained, while the LLM can be LoRA-tuned [22] or fully tuned. First, comparing #0 with #1, or #3 with #4, training the vision abstractor significantly enhances performance. Second, similarly, from #1 to #2, or #4 to #5, training the vision encoder also leads to performance improvements. This improvement may be attributed to the fact that training either the vision encoder or abstractor helps extract more relevant features for IQA. Third, comparing #0 with #3, when the vision components are fixed, fully-tuned LLM shows a substantial advantage over LoRA-tuned LLM. Finally, from #1 to #4, or #2 to #5, when the vision components are trainable, fully-tuned LLM demonstrates similar performance to LoRA-tuned LLM. This provides an alternative if the computation resources are limited.

### v3

@@ 4.2 Score Estimation during Inference
Here we follow Q-Align [qalign] to apply a close-set softmax on the five levels to obtain the probabilities, $p_{i}^{pred}$, to prevent the influence of other textual tokens. Our statistics in Tab. A2 reveal that the trained MLLM will always predict the five levels, thus applying softmax on either the five levels or all textual tokens yields nearly identical results. As given in Tab. 5, our predicted score distribution, $\mathcal{N}(\mu^{pred},(\sigma^{pred})^{2})$, aligns well with human annotations.
@@ 4.3 Joint Training on Multiple IQA Datasets
As demonstrated in [unique, compare2score], different IQA datasets exhibit distinct distributions. Due to this dataset variation, images from different datasets with nearly identical scores (i.e., linearly re-scaled scores) can have drastically different perceptual quality (see examples in Fig. A1). The reason is that different datasets have different reflections between quality scores and visual quality. This disparity makes it infeasible to merely regress absolute quality scores for multi-dataset co-training. To address this challenge, UNIQUE [unique] applies fidelity loss [fidelity] based on Thurstone’s model [thurstone]. The fidelity loss is based on the principal that IQA emphasizes the ranking relationships within the same dataset, rather than the absolute values of quality scores. Specifically, two images are sampled from the same dataset, whose quality scores can be comparable. These two images are then compared using human-annotated quality scores. The fidelity loss encourages the model to learn such comparison results, which facilitates the model to focus more on the ranking relationships in each dataset, instead of the absolute values of quality scores. Thus, the fidelity loss helps multi-dataset co-training. Our soft label is naturally compatible with the fidelity loss, since our method can predict the score distribution, which is what the fidelity loss needs.
@@
Training with fidelity loss. Following [unique], we choose fidelity loss as the similarity measure to train the model as:
@@
Handcrafted NIQE [niqe] 0.533 / 0.530 0.679 / 0.664 0.468 / 0.405 0.195 / 0.161 0.493 / 0.449 0.560 / 0.533 0.718 / 0.628 0.147 / 0.100
BRISQUE [brisque] 0.225 / 0.226 0.490 / 0.406 0.429 / 0.356 0.267 / 0.232 0.361 / 0.313 0.541 / 0.497 0.740 / 0.556 0.108 / 0.054
Non-MLLM Deep-learning NIMA [nima] 0.896 / 0.859 0.838 / 0.856 0.532 / 0.535 0.390 / 0.399 0.814 / 0.771 0.715 / 0.654 0.695 / 0.649 0.561 / 0.467
HyperIQA [HyperIQA] 0.917 / 0.906 0.791 / 0.788 0.506 / 0.468 0.410 / 0.403 0.772 / 0.749 0.702 / 0.640 0.752 / 0.717 0.485 / 0.383
DBCNN [dbcnn] 0.884 / 0.875 0.812 / 0.806 0.497 / 0.484 0.384 / 0.381 0.773 / 0.755 0.730 / 0.641 0.586 / 0.572 0.485 / 0.385
MUSIQ [musiq] 0.924 / 0.929 0.868 / 0.863 0.575 / 0.556 0.431 / 0.431 0.789 / 0.830 0.722 / 0.630 0.771 / 0.710 0.565 / 0.467
CLIP-IQA+ [clipiqa] 0.909 / 0.895 0.866 / 0.864 0.653 / 0.654 0.427 / 0.419 0.832 / 0.805 0.736 / 0.685 0.772 / 0.719 0.427 / 0.316
ManIQA [maniqa] 0.849 / 0.834 0.768 / 0.758 0.499 / 0.465 0.457 / 0.452 0.849 / 0.832 0.723 / 0.636 0.623 / 0.627 0.512 / 0.401
MLLM-based Compare2Score [compare2score] 0.923 / 0.910 0.867 / 0.860 0.500 / 0.453 0.354 / 0.342 0.786 / 0.772 0.777 / 0.671 0.735 / 0.705 0.474 / 0.413
Q-Align [qalign] 0.941 / 0.940 0.886 / 0.887 0.674 / 0.684 0.403 / 0.419 0.853 / 0.860 0.772 / 0.735 0.785 / 0.737 0.554 / 0.483
@@
0 KonIQ, SPAQ Q-Align [qalign] 0.943 / 0.940 0.933 / 0.931 0.692 / 0.708 0.401 / 0.411 0.883 / 0.879 0.795 / 0.727 0.671 / 0.568 0.795 / 0.767
@@
1 KonIQ, SPAQ, KADID Q-Align [qalign] 0.945 / 0.938 0.933 / 0.931 0.935 / 0.934 0.409 / 0.420 0.887 / 0.883 0.788 / 0.733 0.829 / 0.808 0.876 / 0.845
@@
2 KonIQ, SPAQ, KADID, PIPAL Q-Align [qalign] 0.926 / 0.932 0.917 / 0.920 0.950 / 0.954 0.702 / 0.671 0.853 / 0.845 0.765 / 0.722 0.811 / 0.795 0.838 / 0.789
@@ 5.1 Datasets, Baselines, Metrics, and Details
Datasets. For a comprehensive evaluation, we include a wide range of IQA datasets: (a) in-the-wild images: KonIQ [koniq], SPAQ [spaq], LIVE-Wild [livewild], and FLIVE [paq2piq]; (b) synthetic distortions: KADID [kadid], TID2013 [tid2013], and CSIQ [csiq]; (c) model-processed distortions: PIPAL [pipal]; (d) AI-generated images: AGIQA-3K [agiqa]. Among these datasets, KonIQ, SPAQ, KADID, and PIPAL are split into training and test sets. Other datasets are used to assess the out-of-distribution generalization ability. The Mean Opinion Scores (MOS) of these datasets are normalized to the range of [1,5], with the variances normalized accordingly.
Baseline methods. To demonstrate the effectiveness of our distribution-based soft label, we primarily compare with Q-Align [qalign], which utilizes the one-hot label. The base models for DeQA-Score and Q-Align are kept identical for fair comparison. We also include previous IQA methods as baselines, including handcrafted methods: NIQE [niqe] and BRISQUE [brisque], and deep-learning-based methods: NIMA [nima], HyperIQA [HyperIQA], DBCNN [dbcnn], MUSIQ [musiq], CLIP-IQA+ [clipiqa], ManIQA [maniqa], and Compare2Score [compare2score].
Metrics. Following [musiq, qalign, maniqa], we use the Pearson Linear Correlation Coefficient (PLCC) and Spearman Rank-order Correlation Coefficient (SRCC) as metrics to evaluate the score regression performance. Additionally, we employ KL divergence, JS divergence, and Wasserstein distance to assess the results of score distribution prediction.
Implementation details. We use mPLUG-Owl2 [mplugowl2] as our base model, with a CLIP-pretrained ViT-L [clip] as the image encoder, a six-layer Q-Former as the visual abstractor, and LLaMA-2-7B [llama2] as the LLM. The number of visual tokens output by the vision abstractor is set to 64, and the hidden dimension of LLM is 4096. The pretrained weights of mPLUG-Owl2 are used for model initialization. The weighting term $\gamma$ is empirically set to 0.05, adjusting the two loss terms roughly to the same scale. We adopt AdamW [adamw] as the optimizer. The initial learning rate is set to 2e-5 and decays gradually using the cosine decay strategy. Our model is trained with a batch size of 64 for 3 epochs. Using 8 NVIDIA RTX A6000 GPUs, the training process on KonIQ dataset is completed in around 1.5 hours.
@@ 5.2 Results of Score Regression
Score regression results of single-dataset training. We first train our model on a single dataset, i.e., KonIQ, and then evaluate the model on both in-distribution and out-of-distribution datasets, following [attiqa, qalign]. The score regression results are shown in Tab. 3. On the in-distribution KonIQ test set, our model achieves superior score regression results compared to all baseline methods, e.g., a 1.3% improvement in PLCC over Q-Align. Furthermore, on out-of-distribution datasets, our DeQA-Score maintains its advantage with the best overall results across nearly all benchmarks, e.g., a 4.6% improvement in PLCC on LIVE-Wild. These findings confirm the effectiveness of our method.
@@ 5.3 Results of Score Distribution Prediction
Datasets Q-Align [qalign] DeQA-Score (Ours)
@@ 5.4 Ablation Studies
Effectiveness of soft label and fidelity loss is studied and shown in Tab. 6. In #1, “Uncertainty” is utilized to verify the effects of adding annotation variances without considering the score distortion. We directly append uncertainty within the responses, i.e., “The quality of this image is <<level>> with <<uncertainty_level>> uncertainty”. The <<uncertainty_level>> has five possible values: {“minimal”, “low”, “moderate”, “high”, “severe”}, derived by the one-hot discretization. In #2, model trained on one-hot label can not predict a reliable variance, thus we employ binary fidelity loss [liqe], which does not require predicted variance.
@@ B.1 Details of Dataset Construction
According to the $3\sigma$ rule, nearly all samples fall within the range $\mu-3\sigma,\mu+3\sigma]$. Thus, if $3\sigma$ is less than half of the level interval, only the two nearest levels have non-zero probabilities. This criterion provides a way to define the threshold for small variance as $3\sigma\leq d/2$, which simplifies to $\sigma\leq 0.17$. To allow for a slightly relaxed threshold, the threshold for small (normalized) variance is set to $(0.2)^{2}$.
@@ B.2 Details of Methodology
Model architecture. DeQA-Score adopts the architecture of mPLUG-Owl2 [mplugowl2], structured as follows. Specifically, the input images and the question texts are first tokenized, then fused, finally processed by the Large Language Model (LLM) for response generation. (a) Tokenizing input images and texts. We use a pre-trained CLIP pre-trained ViT-L/14 [clip] as the image encoder to convert the input images into visual tokens, with each token having a channel of 1024. The texts are tokenized into textual tokens using the SentencePiece tokenizer [sentencepiece], with each token having a channel of 4096. To bridge the different embedding spaces of visual and textual tokens, a trainable image abstractor, which is a six-layer transformer network, is implemented to map the vision tokens to the hidden dimension of the LLM, which is 4096. The abstractor can also significantly reduce the number of vision tokens to 64, relieving the computing pressure. (b) Token fusion. We integrate the visual tokens into pre-defined positions within the textual tokens as token fusion. (c) Response generation using LLM. The fused tokens are fed into an LLM, which is LLaMA-2-7B [llama2], to generate the final response. The LLM can be either LoRA-tuned [lora] or fully tuned, and their results are similar when the vision components are trainable, as shown in Tab. A7.
@@ B.3 Details of Training and Inference
Training Setup. The pre-trained weights of mPLUG-Owl2 are used for model initialization. The loss weighting term is set to 0.05, bringing the two loss terms to roughly the same scale. We adopt AdamW [adamw] as the optimizer, with an initial learning rate of 2e-5 that gradually decays using a cosine decay strategy. A warmup strategy is applied with a warmup ratio of 0.03. Our model is trained with a batch size of 64 for 3 epochs. Both the vision encoder and abstractor are trainable, and the LLM is fully tuned unless specified. As shown in Tab. A7, LoRA-tuned LLM achieves comparable performance to fully-tuned LLM, providing an alternative for scenarios with limited computation resources. Using 8 RTX A6000 GPUs, training on the KonIQ dataset is completed in about 1.5 hours, while training on the KonIQ, SPAQ, and KADID datasets takes around 4 hours.
Explanation of closed-set softmax during inference. In the main paper, we follow Q-Align [qalign] by applying a closed-set softmax on the five levels to compute the probabilities $p_{i}^{\text{pred}}$, thereby avoiding the influence of other textual tokens. As shown in our statistics in Tab. A2, the probability sum of the five levels after training is very close to 1, indicating that the trained MLLM consistently predicts one of the five levels. Thus, applying softmax over the five levels or across all textual tokens yields nearly identical results.
@@ Appendix C More Results
The results are given in Tab. A3, where the KL divergence of Q-Align is quite large. We explain this as follows. As in Figs. 6, A4, A3 and A5, Q-Align tends to predict a single level token, bringing a quite small predicted variance. That means, the first variance, $\sigma_{1}^{2}$ in Eq. A3, is quite small ($\to 0$), thus the KL divergence goes extremely large ($\to+\infty$).
@@
Results on more AI-generated images. In the main paper, we have included an IQA dataset with AI-generated images, AGIQA-3K [agiqa], for evaluation. Here we further provide evaluation results on three additional AIGC datasets including AIGCIQA2023 [aigciqa2023], AGIN [agin], and AGIQA-1K [agiqa1k] in Tab. A4. Both Q-Align and our DeQA-Score are trained on the KonIQ, SPAQ, and KADID datasets, and then directly evaluated on these unseen AIGC datasets. Our DeQA-Score consistently outperforms the baseline method.
@@
LIQE [liqe] 0.907 / 0.922 0.916 / 0.921 0.928 / 0.929 0.503 / 0.493 0.855 / 0.822 0.689 / 0.650 0.881 / 0.854 0.792 / 0.794
Compare2Score [compare2score] 0.941 / 0.929 0.929 / 0.927 0.952 / 0.949 0.446 / 0.440 0.868 / 0.856 0.787 / 0.733 0.836 / 0.809 0.879 / 0.830
Q-Align [qalign] 0.945 / 0.938 0.933 / 0.931 0.935 / 0.934 0.409 / 0.420 0.887 / 0.883 0.788 / 0.733 0.829 / 0.808 0.876 / 0.845
@@
Table A7: Ablation studies on training / fixing various model components with PLCC / SRCC metrics. “Enc.” means vision encoder, and “Abs.” represents vision abstractor. “LLM LoRA” or “LLM Full” represents the LLM is LoRA-tuned [lora] or fully tuned. “✔” means this component is trained. The models are trained on the KonIQ, SPAQ, and KADID datasets. The results show that training vision encoder and abstractor significantly improves the performance. When vision encoder and abstractor are trained, fully-tuned LLM only shows a slight advantage over LoRA-tuned LLM.
@@
Table A8: Low-level perception results on Q-Bench [qbench] of co-training with or pre-training on score regression tasks. The results of #0, #1, and #2 are borrowed from [qinstruct]. Co-training with or pre-training on score regression tasks stably improves the performance.
@@
0 From the scratch [qinstruct] 0.7218 0.5796 0.5619 0.5668 0.6921 0.5329 0.7265 0.6161
1 High-level co-training [qinstruct] 0.7564 0.6704 0.5903 0.7101 0.6528 0.6316 0.6980 0.6756
2 High-level pre-training [qinstruct] 0.7600 0.6504 0.6166 0.6595 0.6875 0.6546 0.7388 0.6796
@@
Table A9: Co-training with Q-Instruct [qinstruct] leads to an obvious reduction in score regression performance. Exploring better strategies to combine language-based quality description tasks with score regression tasks is left as our future work.
@@
More baselines on multi-dataset training. In the main paper, we primarily compare with Q-Align. Here we add the comparison results with more baselines including LIQE [liqe] and Compare2Score [compare2score] in Tab. A5. Our DeQA-Score consistently outperforms these baselines.
[…]
Ablation studies on training / fixing components are summarized in Tab. A7. The vision encoder and abstractor can be either fixed or trained, while the LLM can be LoRA-tuned [lora] or fully tuned. First, comparing #0 with #1, or #3 with #4, training the vision abstractor significantly enhances performance. Second, similarly, from #1 to #2, or #4 to #5, training the vision encoder also leads to performance improvements. This improvement may be attributed to the fact that training either the vision encoder or abstractor helps extract more relevant features for IQA. Third, comparing #0 with #3, when the vision components are fixed, fully-tuned LLM shows a substantial advantage over LoRA-tuned LLM. Finally, from #1 to #4, or #2 to #5, when the vision components are trainable, fully-tuned LLM demonstrates similar performance to LoRA-tuned LLM. This provides an alternative if the computation resources are limited.